OpenStudy (s117):

Which pair of numbers is relatively prime? A. 21 and 39 B. 10 and 21 C. 12 and 54 D. 15 and 27

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

do you know what "relatively prime" means ?

OpenStudy (s117):

no

OpenStudy (misty1212):

then you can't do it

OpenStudy (misty1212):

you need to know what it means before you can answer the question https://www.mathsisfun.com/definitions/relatively-prime.html

OpenStudy (s117):

@iGreen.

OpenStudy (s117):

@iambatman

OpenStudy (anonymous):

"Two integers are relatively prime if they share no common positive factors."

OpenStudy (anonymous):

21 and 39 2 + 1 = ? 3 + 9 = ?

OpenStudy (s117):

3 12

OpenStudy (anonymous):

Yes, and both of those are divisible by 3, so A is not correct.

OpenStudy (anonymous):

iGreen always beats me to it :), go igreen!

OpenStudy (anonymous):

10 and 21 1 + 0 = ? 2 + 1 = ?

OpenStudy (anonymous):

Thanks :)

OpenStudy (s117):

1 3

OpenStudy (anonymous):

Oh, and FYI, all of these are divisible by 3 except 1 pair, so we just have to find the 1 pair that isn't divisible by 3.

OpenStudy (anonymous):

Okay, are both those numbers divisible by 3?

OpenStudy (s117):

no

OpenStudy (anonymous):

Yes, so B is our answer.

OpenStudy (s117):

thnx @iGreen.